OpenStudy (anonymous):

A flower was found to have two sets of opposing spirals formed by florets. The clockwise spiral has x florets. The number of florets in the counterclockwise direction is 2 less than that in the clockwise direction. These floret numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of florets in the two spirals?

OpenStudy (anonymous):

x by all of x minus 2 is equal to 1.6. x by 2 is equal to 1.6. 2 by x is equal to 1.6. all of x plus 2 by 2 is equal to 1.6.

OpenStudy (anonymous):

i got the 1st one.

jimthompson5910 (jim_thompson5910):

my guess is that the first choice says x/(x-2) = 5/3 which would become x/(x-2) = 1.6 after you divide but I'm not entirely sure

OpenStudy (anonymous):

so you got the same thing i did?

jimthompson5910 (jim_thompson5910):

well if you have x going clockwise and 2 less than that (x-2) going counterclockwise, then we would have x/(x-2) It says that these two numbers are part of the Fibonacci sequence, which is 0, 1, 1, 2, 3, 5, 8, 13, ... So these numbers have to be 3 and 5 since they are the only two that are 2 apart

jimthompson5910 (jim_thompson5910):

So x/(x-2) = 5/3 which becomes x/(x-2) = 1.6 So yes, I'm getting the same thing you are getting

jimthompson5910 (jim_thompson5910):

but this is an odd way to phrase a problem